# Good-for-MDP State Reduction for Stochastic LTL Planning

### v1

@@ 1 Introduction
A major factor affecting the scalability of solving MDPs with LTL goals is the size of the GFM automata corresponding to the LTL goals (Hahn et al. 2015; Sickert and Kretínský 2016). In this paper, we propose a novel GFM state-space reduction technique that significantly reduces the number of automata states. Our method employs a sophisticated chain of transformations, which allows us to leverage recent advances in good-for-games minimisation developed for adversarial settings (Radi and Kupferman 2022) to our context. We provide experimental evidence that our state-reduction technique effectively reduces the automata state space, by taking benchmark examples of LTL goals from over 8 sources, including influential papers in planning, verification and reinforcement learning.
@@ 2 Preliminaries
A sub-MDP of $\mathcal{M}$ is an MDP $\mathcal{M}^{\prime}=(S^{\prime},\text{Act}^{\prime},\Sigma,\mathsf{P}^{\prime},\textit{L})$ where $S^{\prime}\subseteq S,\text{Act}^{\prime}\subseteq\text{Act}$ is such that, for every $s\in S^{\prime}$, $\text{Act}^{\prime}(s)\subseteq\text{Act}(s)$, and $\mathsf{P}^{\prime}$ and $\textit{L}^{\prime}$ are analogous to $\mathsf{P}$ and L when restricted to $S^{\prime}$ and $\text{Act}^{\prime}$. In particular, $\mathcal{M}^{\prime}$ is closed under probabilistic transitions, i.e., for all $s\in S^{\prime}$ and $a\in\text{Act}^{\prime}$ we have that $\mathsf{P}(s,a,s^{\prime})>0$ implies $\mathsf{P}^{\prime}(s,a,s^{\prime})>0$. An end-component (EC) of an MDP $\mathcal{M}$ is a sub-MDP $\mathcal{M}^{\prime}$ of $\mathcal{M}$ such that its underlying graph is strongly connected. A maximal end-component (MEC) is an EC $\mathcal{E}=(E,\text{Act}^{\prime},\Sigma,\mathsf{P}^{\prime},\textit{L})$ such that there is no other EC $\mathcal{E}=(E^{\prime},\text{Act}^{\prime\prime},\Sigma,\mathsf{P}^{\prime\prime},\textit{L})$ such that $E\subset E^{\prime}$. An MEC $E$ that cannot reach states outside $E$ is a leaf component.
@@
Once an end-component $EC$ of an MDP is entered, there is a strategy that visits every state-action combination in $E$ with probability $1$ and stays in $EC$ forever. Moreover, for every strategy the union of the end-components is visited with probability $1$. An infinite path of an MC $\mathcal{M}$ almost surely (with probability $1$) will enter a leaf component.
@@
For an infinite trace $w$, a run of $\mathcal{A}$ on $w$ is an infinite sequence of transitions $\rho=(q_{0},w[0],q_{1})(q_{1},w[1],q_{2})\cdots$ such that, for every $i\geq 0$, $q_{i+1}\in\delta(q_{i},w[i])$. Let $\inf(\rho)$ be the set of transitions that occur infinitely often in a run $\rho$. The acceptance condition $\alpha\subseteq\Delta(\delta)$ is a set of accepting (rejecting, resp.) transitions for Büchi (co-Büchi, resp.). A run $\rho$ satisfies the Büchi (co-Büchi, resp.) acceptance condition $\alpha$ if $\inf(\rho)\cap\alpha\neq\emptyset$ ($\inf(\rho)\cap\alpha=\emptyset$, resp.).
@@ 3 Stochastic Planning Problem
The probability coming from the syntactic construction above is characterised as:
$\textsf{Psyn}(\mathcal{M},\mathcal{A})=\sup_{\mu}\mathbb{P}\{\xi\in\Omega^{\mathcal{M}\times\mathcal{A}}_{\mu}(\langle s_{0},q_{0}\rangle):\xi\text{ accepting}\}$
which can be simplified to $\textsf{Psyn}(\mathcal{M},\mathcal{A})=\sup_{\mu}\mathbb{P}_{(\mathcal{M}\times\mathcal{A})^{\mu}}(\textsf{F}X)$, where $X$ is the set of states of the accepting MECs in $\mathcal{M}\times\mathcal{A}$ that contain accepting transitions in $\alpha^{\times}$.
@@ 4 Probabilistic $\omega$-Automata
Our state reduction technique is based on obtaining a small probabilistic (Büchi) automata (PAs) out of a GFM automaton. The key characteristic of PAs (Baier, Größer, and Bertrand 2012) is that, while being nondeterministic, their nondeterministic choices are resolved randomly. A PA $\mathcal{P}=(\Sigma,Q,\delta,q_{0},\alpha)$ is a nondeterministic automaton equipped with a Büchi acceptance condition $\alpha$ and a randomised transition function $\delta:Q\times\Sigma\mapsto\mathsf{Distr}(Q)$, where, from state $q$ and letter $\sigma$, transition to $q^{\prime}$ is taken with probability $\delta(q,\sigma)(q^{\prime})$. We often abuse notation by writing $\delta(q,\sigma)$ to denote its support. Each word $w\in\Sigma^{\omega}$ induces a probability measure $\mathbb{P}^{w}_{\mathcal{P}}$ on $Q^{\omega}$ in the usual way. The probability that $\mathcal{P}$ accepts $w$, denoted by $\mathbb{P}_{\mathcal{P}}(w)$, is the probability measure of all accepting runs of $w$ on $\mathcal{P}$, that is:
$\mathbb{P}_{\mathcal{P}}(w)=\mathbb{P}^{w}_{\mathcal{P}}(\{\pi:\pi\text{ is an accepting run of }w\})$
A PA $\mathcal{P}$ is 0/1-PA if, for any word $w\in\Sigma^{\omega}$, we have either $\mathbb{P}_{\mathcal{P}}(w)=1$ or $\mathbb{P}_{\mathcal{P}}(w)=0$. For a 0/1-PA $\mathcal{P}$, with a slight abuse of notation, we say a word $w$ is accepted by $\mathcal{P}$ ($w$ is in $\mathcal{L}(\mathcal{P})$) if $\mathbb{P}_{\mathcal{P}}(w)=1$. 0/1-PAs are known to be semantically deterministic (Li et al. 2025). That is, given a state $p$ and a letter $\sigma$ of 0/1-PA $\mathcal{P}$, for any two states $q,r\in\delta(p,\sigma)$, we have that $\mathcal{L}(\mathcal{P}^{q})=\mathcal{L}(\mathcal{P}^{r})$ where $\mathcal{P}^{s}$ is the automaton by setting the initial state to $s$. For a detailed introduction to PAs please refer to (Baier, Größer, and Bertrand 2012).
@@
In the following, we show for the first time that 0/1-PAs can also be perceived as another type of GFM automata. To this end, we first define the product MDP of an MDP $\mathcal{M}$ and a 0/1-PA $\mathcal{P}$  [^3] We note that a source of 0/1-PAs can be a special type of automata whose nondeterminism can be resolved by randomness discussed in (Henzinger, Prakash, and Thejaswini 2025). Such kind of automata is also said to be GFM automata in a discussion without formal proofs in (Henzinger, Prakash, and Thejaswini 2025)..
@@
  • $\alpha^{\otimes}=\{(\langle s,q\rangle,a,\langle s^{\prime},q^{\prime}\rangle)\mid(q,\sigma,q^{\prime})\in\alpha,a\in\text{Act}^{\otimes},\mathsf{P}^{\otimes}(\langle s,q\rangle,a,\langle s^{\prime},q^{\prime}\rangle)>0,\sigma=\textit{L}^{\otimes}(\langle s,q\rangle,a)\}$.
@@ 5 Good-for-MDP State Reduction
In this section, we present our main contribution, a general state space reduction for GFM automata. We assume to have a GFM automaton $\mathcal{A}_{\text{GFM}}$ and through a polynomial Redux procedure, we return a 0/1-PA $\mathcal{A}_{\text{PA}}$, whose state-space can be significantly reduced. Redux works in several stages, see Figure 1, which we detail below.
@@
  • $\textit{L}^{\prime}:S^{\prime}\times\text{Act}^{\prime}\rightarrow\Sigma^{\prime}$ where if $\textit{L}(s,a)=\sigma$, then $\textit{L}^{\prime}(s,\langle a,i\rangle)=\langle\sigma,i\rangle$ for all $i\in[k]$, and
@@
GFG automata (Henzinger and Piterman 2006) are automata in which nondeterminism can be resolved deterministically by assigning the choice to the protagonist in an adversarial (vs. stochastic MDPs) games. Formally, an automaton $\mathcal{A}$ is said to be GFG if there exists a strategy $f:\Sigma^{*}\rightarrow Q$ such that for every word $w=\sigma_{0}\sigma_{1}\cdots$, the run $f(\epsilon)f(\sigma_{0})\cdots f(\sigma_{0}\cdots\sigma_{i})\cdots$ is accepting. Intuitively, this indicates that there is a deterministic strategy to produce an accepting run in a GFG automaton even if the input word is given letter by letter. Interestingly, GFG co-Büchi automata can be minimised in polynomial time (Radi and Kupferman 2022). Since GFG automata have more restricted nondeterminism than GFM automata, GFG automata are also GFM (Klein et al. 2014). GFG automata are not adopted in state-of-the-art stochastic planning tools, such as PRISM (Kwiatkowska, Norman, and Parker 2011). The reason is that GFG Büchi (resp. co-Büchi) automata cannot recognise all languages expressed by LTL as they are only as expressive as their deterministic counterparts. Moreover, current approaches to construct GFG automata are not as efficient as the ones for GFM. Notably, our work is able to use GFG minimisation on GFM automata that recognise all $\omega$-regular properties. See Appendix G for more details.
@@
  • Identify the accepting MECs that have some accepting transition.
  • Synthesise an optimal strategy $\mu$ that maximises the reachability probability of accepting MECs in usual way.
@@ 6 Direct GFM Construction for $\textsf{G}\textsf{F}\varphi$
The $\textsf{G}\textsf{F}\varphi$ fragment is very common in verification (Holeček et al. 2004) and also in reinforcement learning (Jackermeier and Abate 2025). More generally, repeated reachability properties correspond to a commonly used subclass of so-called recurrence properties (Manna and Pnueli 1990; Chang, Manna, and Pnueli 1992).
@@ 7 Experiments for GFM state reduction
In this section, we validate the performance of our GFM state-space reduction on LTL specifications extracted from literature, which we define blow. The full benchmark set, demonstrating consistent state space reduction on complex LTL specifications from over 8 sources, is in Appendix N.
@@ 8 Experiments for $\textsf{G}\textsf{F}\varphi$
Table 2 compares our direct GFM construction (denoted by GFM-GF) to the automata obtained by our GFM state-space reduction. Again, we denote number of states of reduced Owl and “slim” GFM automata (Owl-Red, Slim-Red) but in contrast to Table 1, the runtime (in seconds), now contains both LTL to GFM construction and subsequent application of our reduction.

### v2

@@ 1 Introduction
A major factor affecting the scalability of solving MDPs with LTL goals is the size of the GFM automata corresponding to the LTL goals (Hahn et al. 2015; Sickert and Kretínský 2016). In this paper, we propose a novel GFM state-space reduction technique that significantly reduces the number of automata states. Our method employs a sophisticated chain of transformations, which allows us to leverage recent advances in good-for-games minimisation developed for adversarial settings (Radi and Kupferman 2022) to our context. We provide experimental evidence that our state-reduction technique effectively reduces the automata state space, by taking benchmark examples of LTL goals from over eight sources, including influential papers in planning, verification and reinforcement learning.
@@ 2 Preliminaries
A sub-MDP of $\mathcal{M}$ is an MDP $\mathcal{M}^{\prime}=(S^{\prime},\text{Act}^{\prime},\Sigma,\mathsf{P}^{\prime},\textit{L})$ where $S^{\prime}\subseteq S,\text{Act}^{\prime}\subseteq\text{Act}$ is such that, for every $s\in S^{\prime}$, $\text{Act}^{\prime}(s)\subseteq\text{Act}(s)$, and $\mathsf{P}^{\prime}$ and $\textit{L}^{\prime}$ are analogous to $\mathsf{P}$ and L when restricted to $S^{\prime}$ and $\text{Act}^{\prime}$. In particular, $\mathcal{M}^{\prime}$ is closed under probabilistic transitions, i.e., for all $s\in S^{\prime}$ and $a\in\text{Act}^{\prime}$ we have that $\mathsf{P}(s,a,s^{\prime})>0$ implies $\mathsf{P}^{\prime}(s,a,s^{\prime})>0$. An end-component (EC) of an MDP $\mathcal{M}$ is a sub-MDP $\mathcal{M}^{\prime}$ of $\mathcal{M}$ such that its underlying graph is strongly connected. A maximal end-component (MEC) is an EC $\mathcal{E}=(E,\text{Act}^{\prime},\Sigma,\mathsf{P}^{\prime},\textit{L})$ such that there is no other EC $\mathcal{E}=(E^{\prime},\text{Act}^{\prime\prime},\Sigma,\mathsf{P}^{\prime\prime},\textit{L})$ such that $E\subset E^{\prime}$. An MEC $\mathcal{E}$ that cannot reach states outside $\mathcal{E}$ is a leaf component.
@@
Once an EC $\mathcal{E}$ of an MDP is entered, there is a strategy that visits every state-action combination in $\mathcal{E}$ with probability $1$ and stays in $\mathcal{E}$ forever. Moreover, for every strategy the union of the end-components is visited with probability $1$. An infinite path of an MC $\mathcal{M}$ almost surely (with probability $1$) will enter a leaf component.
@@
For an infinite trace $w\in\Sigma^{\omega}$, a run of $\mathcal{A}$ on $w$ is an infinite sequence of transitions $\rho=(q_{0},w[0],q_{1})(q_{1},w[1],q_{2})\cdots$ such that, for every $i\geq 0$, $q_{i+1}\in\delta(q_{i},w[i])$. Let $\inf(\rho)$ be the set of transitions that occur infinitely often in a run $\rho$. The acceptance condition $\alpha\subseteq\Delta(\delta)$ is a set of accepting (rejecting, resp.) transitions for Büchi (co-Büchi, resp.). A run $\rho$ satisfies the Büchi (co-Büchi, resp.) acceptance condition $\alpha$ if $\inf(\rho)\cap\alpha\neq\emptyset$ ($\inf(\rho)\cap\alpha=\emptyset$, resp.).
@@ 3 Stochastic Planning Problem
The probability coming from the syntactic product construction above is characterised as:
$\textsf{Psyn}(\mathcal{M},\mathcal{A})=\sup_{\mu}\mathbb{P}\{\xi\in\Omega^{\mathcal{M}\times\mathcal{A}}_{\mu}(\langle s_{0},q_{0}\rangle):\xi\text{ accepting}\}.$
This can be simplified to $\textsf{Psyn}(\mathcal{M},\mathcal{A})=\sup_{\mu}\mathbb{P}_{(\mathcal{M}\times\mathcal{A})^{\mu}}(\textsf{F}X)$, where $X$ is the set of states of the accepting MECs in $\mathcal{M}\times\mathcal{A}$ that contain accepting transitions in $\alpha^{\times}$.
@@ 4 Probabilistic $\omega$-Automata
Our state reduction technique is based on obtaining a small probabilistic (Büchi) automaton (PA) out of a GFM automaton. The key characteristic of PAs (Baier, Größer, and Bertrand 2012) is that, while being nondeterministic, their nondeterministic choices are resolved randomly. A PA $\mathcal{P}=(\Sigma,Q,\delta,q_{0},\alpha)$ is a nondeterministic automaton equipped with a Büchi acceptance condition $\alpha$ and a randomised transition function $\delta:Q\times\Sigma\mapsto\mathsf{Distr}(Q)$, where, from state $q$ and letter $\sigma$, transition to $q^{\prime}$ is taken with probability $\delta(q,\sigma)(q^{\prime})$. We often abuse notation by writing $\delta(q,\sigma)$ to denote its support. Each word $w\in\Sigma^{\omega}$ induces a probability measure $\mathbb{P}^{w}_{\mathcal{P}}$ on $Q^{\omega}$ in the usual way. The probability that $\mathcal{P}$ accepts $w$, denoted by $\mathbb{P}_{\mathcal{P}}(w)$, is the probability measure of all accepting runs of $w$ on $\mathcal{P}$, that is:
$\mathbb{P}_{\mathcal{P}}(w)=\mathbb{P}^{w}_{\mathcal{P}}(\{\pi:\pi\text{ is an accepting run of }w\}).$
A PA $\mathcal{P}$ is a 0/1-PA if, for any word $w\in\Sigma^{\omega}$, we have either $\mathbb{P}_{\mathcal{P}}(w)=1$ or $\mathbb{P}_{\mathcal{P}}(w)=0$. For a 0/1-PA $\mathcal{P}$, with a slight abuse of notation, we say a word $w$ is accepted by $\mathcal{P}$ ($w$ is in $\mathcal{L}(\mathcal{P})$) if $\mathbb{P}_{\mathcal{P}}(w)=1$. 0/1-PAs are known to be semantically deterministic (Li et al. 2025). That is, given a state $p$ and a letter $\sigma$ of 0/1-PA $\mathcal{P}$, for any two states $q,r\in\delta(p,\sigma)$, we have that $\mathcal{L}(\mathcal{P}^{q})=\mathcal{L}(\mathcal{P}^{r})$ where $\mathcal{P}^{s}$ is the automaton by setting the initial state to $s$. For a detailed introduction to PAs please refer to (Baier, Größer, and Bertrand 2012).
@@
In the following, we show for the first time that 0/1-PAs can also be perceived as another type of GFM automata. To this end, we first define the product MDP of an MDP $\mathcal{M}$ and a 0/1-PA $\mathcal{P}$  [^3] We note that a source of 0/1-PAs can be a special type of automata whose nondeterminism can be resolved by randomness discussed in (Henzinger, Prakash, and Thejaswini 2025). This type of automata is also said to be GFM in a discussion without formal proofs in (Henzinger, Prakash, and Thejaswini 2025)..
@@
  • $\alpha^{\otimes}=\{(\langle s,q\rangle,a,\langle s^{\prime},q^{\prime}\rangle)\mid(q,L(s,a),q^{\prime})\in\alpha,(q,\sigma,q^{\prime})\in\alpha,a\in\text{Act}^{\otimes},\mathsf{P}^{\otimes}(\langle s,q\rangle,a,\langle s^{\prime},q^{\prime}\rangle)>0\}$.
@@ 5 Good-for-MDP State Reduction
In this section, we present our main contribution, a general state space reduction for GFM automata. We assume to have a GFM automaton $\mathcal{A}_{\text{GFM}}$ and through a polynomial Redux procedure, we return a 0/1-PA $\mathcal{A}_{\text{PA}}$, whose state space can be significantly reduced. Redux works in several stages, see Figure 1, which we detail below.
@@
  • $\textit{L}^{\prime}:S^{\prime}\times\text{Act}^{\prime}\rightarrow\Sigma^{\prime}$ where $\Sigma^{\prime}=\Sigma\times[k]$, and if $\textit{L}(s,a)=\sigma$, then $\textit{L}^{\prime}(s,\langle a,i\rangle)=\langle\sigma,i\rangle$ for all $i\in[k]$, and
@@
GFG automata (Henzinger and Piterman 2006) are automata in which nondeterminism can be resolved deterministically by assigning the choice to the protagonist in an adversarial (vs. stochastic MDPs) games. Formally, an automaton $\mathcal{A}$ is said to be GFG if there exists a strategy $f:\Sigma^{*}\rightarrow Q$ such that for every accepting word $w=\sigma_{0}\sigma_{1}\cdots$, the run $f(\epsilon)f(\sigma_{0})\cdots f(\sigma_{0}\cdots\sigma_{i})\cdots$ is accepting. Intuitively, this indicates that there is a deterministic strategy to produce an accepting run in a GFG automaton even if the input accepting word is given letter by letter. Interestingly, GFG co-Büchi automata can be minimised in polynomial time (Radi and Kupferman 2022). Since GFG automata have more restricted nondeterminism than GFM automata, GFG automata are also GFM (Klein et al. 2014). GFG automata are not adopted in state-of-the-art stochastic planning tools, such as PRISM (Kwiatkowska, Norman, and Parker 2011). The reason is that GFG Büchi (resp. co-Büchi) automata cannot recognise all languages expressed by LTL as they are only as expressive as their deterministic counterparts. Moreover, current approaches to construct GFG automata are not as efficient as the ones for GFM. Notably, our work is able to use GFG minimisation on GFM automata that recognise all $\omega$-regular properties. See Appendix G for more details.
@@
  • Identify the accepting MECs that have some accepting transitions.
  • Synthesise an optimal strategy $\mu$ that maximises the reachability probability of accepting MECs as usual.
@@ 6 Direct GFM Construction for $\textsf{G}\textsf{F}\varphi$
The $\textsf{G}\textsf{F}\varphi$ fragment is very common in verification (Holeček et al. 2004) and also in reinforcement learning (Jackermeier and Abate 2025). More generally, repeated reachability properties are within a commonly used subclass of so-called recurrence properties (Manna and Pnueli 1990; Chang, Manna, and Pnueli 1992).
@@ 7 Experiments for GFM state reduction
In this section, we validate the performance of our GFM state space reduction on LTL specifications extracted from literature, which we define blow. The full benchmark set, demonstrating consistent state space reduction on complex LTL specifications from over 8 sources, is in Appendix N.
@@ 8 Experiments for $\textsf{G}\textsf{F}\varphi$
Table 2 compares our direct GFM construction (denoted by GFM-GF) to the automata obtained by our GFM state space reduction. Again, we denote number of states of reduced Owl and “slim” GFM automata (Owl-Red, Slim-Red) but in contrast to Table 1, the runtime (in seconds), now contains both LTL to GFM construction and subsequent application of our reduction.